# literature.bib

### v2

@@ 3.3 Convergence to Overdamped Langevin Diffusion
$\qquad+\varepsilon\int\left[(\tilde{y}(\tilde{x})-y)\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)+\lVert M(\tilde{x})\rVert_{F}^{2}\right]\,\mathrm{d}\Pi_{t}$
@@
$\int(y-\tilde{y}(\tilde{x}))\cdot(q-K\tilde{x})\,\mathrm{d}\Pi_{t}\geq\frac{\omega_{f^{\ast}}}{\varepsilon}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\varepsilon}^{2}\right]$
@@
$\varepsilon\int\lVert M(\tilde{x})\rVert_{F}^{2}\,\mathrm{d}\Pi_{t}\leq\varepsilon C_{1}.$
@@
$\varepsilon\int(\tilde{y}(\tilde{x})-y)$ $\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)\,\mathrm{d}\Pi_{t}$
$\leq\varepsilon\int\lVert\tilde{y}(\tilde{x})-y\rVert\cdot\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert\,\mathrm{d}\Pi_{t}$
@@
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\varepsilon}^{2}\right]+\frac{\varepsilon}{2\delta}\mathbb{E}_{\tilde{x}\sim\mu^{\ast}}\left[\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert^{2}\right]$
@@
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\varepsilon}^{2}\right]+\frac{\varepsilon(C_{2}+C_{3})}{2\delta}.$
@@
Denoting $\varphi(t):=\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\varepsilon}(x,y,\tilde{x})^{2}\right]$, we therefore have
$\varphi^{\prime}(t)\leq-2\omega\varphi(t)+2\varepsilon C_{1}+\frac{\varepsilon(C_{2}+C_{3})}{\delta},$
@@
where we abbreviate $\omega\coloneqq\min\{\omega_{g},\frac{\omega_{f^{\ast}}}{\varepsilon}-\frac{\delta}{2}\}$. Using Gronwall’s lemma, this implies
$\varphi(t)$ $\leq e^{-2\omega t}\varphi(0)+\int_{0}^{t}e^{-2\omega(t-s)}\left(2\varepsilon C_{1}+\frac{\varepsilon(C_{2}+C_{3})}{\delta}\right)\,\mathrm{d}s$
$=e^{-2\omega t}\varphi(0)+\left(2\varepsilon C_{1}+\frac{\varepsilon(C_{2}+C_{3})}{\delta}\right)\frac{1-e^{-2\omega t}}{2\omega}$
@@
By the definition of $\mathcal{T}_{1,\varepsilon}$ and taking the infimum over all possible initializations $\Pi_{0}$ with $(x,y)$-marginal $\pi_{0}$ and $(\tilde{x},\tilde{y})$-marginal $\pi^{\ast}=(id,\nabla f\circ K)_{\#}\mu^{\ast}$ on the right side, we get
$\mathcal{T}_{1,\varepsilon}^{2}(\pi_{t},\pi^{\ast})\leq\varphi(t)\leq e^{-2\omega t}\mathcal{T}_{1,\varepsilon}^{2}(\pi_{0},\pi^{\ast})+\left(2\varepsilon C_{1}+\frac{\varepsilon(C_{2}+C_{3})}{\delta}\right)\frac{1-e^{-2\omega t}}{2\omega}.$ (3.12)
@@
For $\varepsilon<\omega_{f^{\ast}}/\omega_{g}$, the bound on the right is minimized by $\delta=2(\frac{\omega_{f^{\ast}}}{\varepsilon}-\omega_{g})$ for any finite $t$ as well as in the limit $t\to\infty$. This choice implies $\omega=\omega_{g}$ and we denote the resulting constant terms by
$C(\varepsilon)\coloneqq\left(\frac{\varepsilon C_{1}}{\omega_{g}}+\frac{\varepsilon(C_{2}+C_{3})}{4\omega_{g}(\omega_{f^{\ast}}-\varepsilon\omega_{g})}\right)^{1/2}=\sqrt{\varepsilon C_{1}/\omega_{g}}+\mathcal{O}(\varepsilon^{3/2}).$
@@
The first desired inequality for $t>0$ follows by taking the square root of (3.12) and using that $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\mathcal{T}_{1,\varepsilon}(\pi_{t},\pi^{\ast})$ and, assuming a concentrated initialization $Y_{0}=\nabla f(KX_{0})$, that
$\mathcal{T}_{1,\varepsilon}(\pi_{0},\pi^{\ast})\leq(1+\varepsilon\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast}).$
@@
Since $\pi_{t}$ converges to the stationary distribution of interest $\pi^{\text{{stat}}}_{\text{{cPD}}}$, taking the limit $t\to\infty$ in (3.12) and using $\mathcal{W}_{2}(\mu^{\text{{stat}}}_{\text{{cPD}}},\mu^{\ast})\leq\mathcal{T}_{1,\varepsilon}(\pi^{\text{{stat}}}_{\text{{cPD}}},\pi^{\ast})$ proves the second statement. ∎
@@ 3.4 A Bias-Corrected Modification of Primal-Dual Diffusion
Note that $B_{\text{{mod}}}$ differs from $B_{\text{{cPD}}}$ in a non-homogeneous coefficient in the dual part $y$. The drift $A_{\text{{mod}}}$ contains the same terms as $A_{\text{{cPD}}}$, and an additional term $-M(x)^{T}\nabla\log\mu^{\ast}(x)$.
@@
In principle, this theorem might suggest that a time discretization of the modified (3.13) may be a better idea than the primal-dual sampling algorithm (2.4). Note, however, apart from the higher-order derivative terms in $f$ (which counter the motivation to use a primal-dual scheme in the first place), the term $A_{\text{{mod}}}$ contains just the gradient of the target log-density, i.e. we explicitly add overdamped Langevin diffusion in the dual variable. This suggests that algorithmically, the dualization is somewhat redundant and any sampler would be inferior to a standard unadjusted Langevin algorithm. In order to prove Theorem 3.13, we need the following vector calculus identity.
@@
$\int\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}^{T}\begin{pmatrix}I&M(x)\\ M(x)^{T}&M(x)M(x)^{T}\end{pmatrix}\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}\varphi\,\mathrm{d}\pi^{\ast}-\int\nabla_{x}\varphi\cdot(\nabla g(x)+K^{*}y)\,\mathrm{d}\pi^{\ast}$ (3.16a)
@@
$-\int\nabla_{x}$ $\varphi\cdot(\nabla g(x)+K^{*}y)\,\mathrm{d}\pi^{\ast}$
$=\int\nabla_{x}\varphi(x,\nabla f(Kx))\cdot(-\nabla g(x)-K^{*}\nabla f(Kx))\exp(-f(Kx)-g(x))\,\mathrm{d}x$
@@
In general, it is hard to generalize a contraction result like Lemma 3.3 to the case of non-homogeneous diffusion coefficients. In section 2 of [Fournier2021], this question is addressed in more detail for the heat equation with non-homogeneous diffusion. The authors show that it becomes possible to prove stability in a Wasserstein distance with a modified underlying metric, where the metric depends on the diffusion coefficient via an elliptic PDE.
@@
4 Stability Analysis in Discrete Time
We want to apply similar techniques as in Section 3 to the time discretized setting of the algorithm (2.4). Indeed, couplings like the ones we employed in the continuous time setting above have been used before in the stability and convergence analysis of Langevin algorithms. We begin this section by recalling such a computation in the simple case of ULA as an instructive example, and then expand to the primal-dual algorithm considered here.
Consider two Markov chains $(X_{n})_{n},(\tilde{X}_{n})_{n}$ defined by initial values $X_{0},\tilde{X}_{0}$ and
$\xi^{n+1}$ $\sim\mathrm{N}(0,I)$
$X^{n+1}$ $=X^{n}-\tau\nabla h(X_{n})+\sqrt{2\tau}\xi^{n+1}$
$\tilde{X}^{n+1}$ $=\tilde{X}^{n}-\tau\nabla h(\tilde{X}^{n})+\sqrt{2\tau}\xi^{n+1}$
@@ 4 Stability Analysis in Discrete Time
We can use the strong convexity of $h$, which guarantees
$D_{h,\textrm{sym}}(X^{n},\tilde{X}^{n})\geq\frac{\mu_{h}L}{\mu_{h}+L}\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\frac{1}{\mu_{h}+L}\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2},$
@@
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2}\leq\left(1-\tau\frac{2\mu_{h}L}{\mu_{h}+L}\right)\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\tau\left(\tau-\frac{2}{\mu_{h}+L}\right)\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2}.$
If the step size is chosen to be $\tau=2/(\mu_{h}+L)$, we obtain the convergence rate
$\lVert X^{n}-\tilde{X}^{n}\rVert^{2}\leq\left(\frac{L-\mu_{h}}{L+\mu_{h}}\right)^{2n}\lVert X^{0}-\tilde{X}^{0}\rVert^{2}.$
Denoting the distribution of $X^{n}$ by $\mu^{n}$ and that of $\tilde{X}^{n}$ by $\tilde{\mu}^{n}$, we obtain by taking the infimum over all couplings of $\mu^{0}$ and $\tilde{\mu}^{0}$ on the right side that
$\mathcal{W}_{2}(\mu_{n},\tilde{\mu}_{n})\leq\left(\frac{L-\mu_{h}}{L+\mu_{h}}\right)^{n}\mathcal{W}_{2}(\mu_{0},\tilde{\mu}_{0}).$
This shows that ULA is stable and will converge to its unique stationary distribution. Similar computations can be carried out for other time-discretized schemes based on Langevin diffusion, e.g. time-implicit leading to proximal sampling schemes. Since the stationary distribution of the time-discrete scheme is not the target distribution $\mu^{\ast}$, the missing step to showing convergence of the Markov chain is to couple a time-discrete chain and a time-continuous process and then finding an error bound. Typically, this is more challenging and the introduced bias (in Wasserstein distance) between the two settings scales as $\sqrt{\tau}$.
We now turn to the primal-dual setting of (2.4). Consider again two Markov chains with potentially different initial values $(X^{0},Y^{0})\sim\pi^{0}$ and $(\tilde{X}^{0},\tilde{Y}^{0})\sim\tilde{\pi}^{0}$, but equal stochastic input:
$\begin{array}[]{lcl}\xi^{n+1}\sim\mathrm{N}(0,I)&&\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(Y^{n}+\sigma KX_{\theta}^{n})&&\tilde{Y}^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(\tilde{Y}^{n}+\sigma K\tilde{X}_{\theta}^{n})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}&&\tilde{X}^{n+1}=\operatorname{prox}_{\tau g}(\tilde{X}^{n}-\tau K^{T}\tilde{Y}^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})&&\tilde{X}_{\theta}^{n+1}=\tilde{X}^{n+1}+\theta(\tilde{X}^{n+1}-\tilde{X}^{n})\end{array}$ (4.1)

### v3

@@ 3.3 Convergence to Overdamped Langevin Diffusion
$\qquad+\frac{1}{\lambda}\int\left[(\tilde{y}(\tilde{x})-y)\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)+\lVert M(\tilde{x})\rVert_{F}^{2}\right]\,\mathrm{d}\Pi_{t}$
@@
$\int(y-\tilde{y}(\tilde{x}))\cdot(q-K\tilde{x})\,\mathrm{d}\Pi_{t}\geq\lambda\omega_{f^{\ast}}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]$
@@
$\lambda^{-1}\int\lVert M(\tilde{x})\rVert_{F}^{2}\,\mathrm{d}\Pi_{t}\leq\lambda^{-1}D_{1}.$
@@
$\lambda^{-1}\int(\tilde{y}(\tilde{x})-y)$ $\cdot\left(m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\right)\,\mathrm{d}\Pi_{t}$
$\leq\lambda^{-1}\int\lVert\tilde{y}(\tilde{x})-y\rVert\cdot\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert\,\mathrm{d}\Pi_{t}$
@@
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]+\frac{1}{2\delta\lambda}\mathbb{E}_{\tilde{x}\sim\mu^{\ast}}\left[\lVert m(\tilde{x})-M^{T}(\tilde{x})(\tilde{p}+K^{T}\tilde{y}(\tilde{x}))\rVert^{2}\right]$
@@
$\leq\frac{\delta}{2}\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[\lVert y-\tilde{y}(\tilde{x})\rVert_{\lambda^{-1}}^{2}\right]+\frac{D_{2}+D_{3}}{2\delta\lambda}.$
@@
Denoting $\varphi(t):=\mathbb{E}_{(x,y,\tilde{x})\sim\Pi_{t}}\left[d_{1,\lambda^{-1}}(x,y,\tilde{x})^{2}\right]$, we therefore have
$\varphi^{\prime}(t)\leq-2\omega\varphi(t)+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta),$
@@
where we denoted $\omega\coloneqq\omega(\delta)\coloneqq\min\{\omega_{g},\lambda\omega_{f^{\ast}}-\frac{\delta}{2}\}$. Using Gronwall’s lemma, this implies
$\varphi(t)$ $\leq e^{-2\omega t}\varphi(0)+\int_{0}^{t}e^{-2\omega(t-s)}\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\,\mathrm{d}s$
$=e^{-2\omega t}\varphi(0)+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\frac{1-e^{-2\omega t}}{2\omega}$
@@
By the definition of $\mathcal{T}_{1,\lambda^{-1}}$ and taking the infimum over all possible initializations $\Pi_{0}$ with $(x,y)$-marginal $\pi_{0}$ and $(\tilde{x},\tilde{y})$-marginal $\pi^{\ast}=(id,\nabla f\circ K)_{\#}\mu^{\ast}$ on the right side, we get
$\mathcal{T}_{1,\lambda^{-1}}^{2}(\pi_{t},\pi^{\ast})\leq\varphi(t)\leq e^{-2\omega t}\mathcal{T}_{1,\lambda^{-1}}^{2}(\pi_{0},\pi^{\ast})+\lambda^{-1}(2D_{1}+(D_{2}+D_{3})/\delta)\frac{1-e^{-2\omega t}}{2\omega}.$ (3.11)
@@
Bearing in mind the dependency of $\omega$ on $\delta$, for $\lambda<\omega_{g}/\omega_{f^{\ast}}$ the bound on the right is minimized by $\delta=2(\lambda\omega_{f^{\ast}}-\omega_{g})$ for any finite $t$ as well as in the limit $t\to\infty$. This choice implies $\omega=\omega_{g}$ and we denote the resulting constant terms by
$C(\lambda)\coloneqq\left(\frac{D_{1}}{\lambda\omega_{g}}+\frac{D_{2}+D_{3}}{4\omega_{g}(\lambda\omega_{f^{\ast}}-\omega_{g})}\right)^{1/2}=(\lambda\omega_{g}/D_{1})^{-1/2}+\mathcal{O}(\lambda^{-3/2}).$
@@
The first desired inequality for $t>0$ follows by taking the square root of (3.11) and using that $\mathcal{W}_{2}(\mu_{t},\mu^{\ast})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{t},\pi^{\ast})$ and, assuming a concentrated initialization $Y_{0}=\nabla f(KX_{0})$, that
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\pi^{\ast})\leq(1+\lambda^{-1}\omega_{f^{\ast}}\lVert K\rVert)^{1/2}\mathcal{W}_{2}(\mu_{0},\mu^{\ast}).$
@@
Since $\pi_{t}$ converges to the stationary distribution $\pi_{\lambda}$, taking the limit $t\to\infty$ in (3.11) and using $\mathcal{W}_{2}(\mu_{\lambda},\mu^{\ast})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{\lambda},\pi^{\ast})$ proves the second statement. ∎
@@ 3.4 A Bias-Corrected Modification of Primal-Dual Diffusion
Note that $B_{\text{{mod}}}$ depends on $x$ and is therefore not a special case of $B_{\text{{gPD}}}$. The drift $A_{\text{{mod}}}$ contains the same terms as $A_{\text{{PD}}}$, and an additional term $-M(x)^{T}\nabla\log\mu^{\ast}(x)$.
@@
In principle, this theorem might suggest that a time discretization of the modified (3.12) may be a better idea than the primal-dual sampling algorithm (1.4). Note, however, apart from the higher-order derivative terms in $f$ (which counter the motivation to use a primal-dual scheme in the first place), the term $A_{\text{{mod}}}$ contains just the gradient of the target log-density, i.e. we explicitly add overdamped Langevin diffusion in the dual variable. This suggests that, algorithmically, the dualization is somewhat redundant and any sampler based on (3.12) would be inferior to a standard unadjusted Langevin algorithm.
In order to prove Theorem 3.13, we need the following vector calculus identity.
@@
$\int\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}^{T}\begin{pmatrix}I&M(x)\\ M(x)^{T}&M(x)M(x)^{T}\end{pmatrix}\begin{pmatrix}\nabla_{x}\\ \nabla_{y}\end{pmatrix}\varphi\,\mathrm{d}\pi^{\ast}-\int\nabla_{x}\varphi\cdot(\nabla g(x)+K^{T}y)\,\mathrm{d}\pi^{\ast}$ (3.15a)
@@
$-\int\nabla_{x}$ $\varphi\cdot(\nabla g(x)+K^{T}y)\,\mathrm{d}\pi^{\ast}$
$=\int\nabla_{x}\varphi(x,\nabla f(Kx))\cdot(-\nabla g(x)-K^{T}\nabla f(Kx))\exp(-f(Kx)-g(x))\,\mathrm{d}x$
@@
While the the last result showed that the target is now the invariant solution, the asymptotic stability of the new diffusion process is unclear. In general, it is hard to generalize a contraction result like Lemma 3.3 to the case of non-homogeneous diffusion coefficients. In section 2 of [Fournier2021], this question is addressed in more detail for the heat equation with non-homogeneous diffusion. The authors show that it becomes possible to prove stability in a Wasserstein distance with a modified underlying metric, where the metric depends on the diffusion coefficient via an elliptic PDE.
@@
4 Convergence of the Discrete Time Algorithm
We want to apply similar techniques as in Section 3 to the time discretized setting of the algorithm (1.4). Indeed, couplings like the ones we employed in the continuous time setting above have been used before in the stability and convergence analysis of Langevin algorithms. We begin this section by recalling such a computation in the simple case of ULA as an instructive example, and then expand to the primal-dual algorithm.
For a convex functional $h:\mathcal{X}\to\mathbb{R}\cup\{\infty\}$, points $x,\tilde{x}\in\mathcal{X}$ and subgradients $p\in\partial h(x),\tilde{p}\in\partial h(\tilde{x})$, we will use the symmetric Bregman distance $D_{h,\text{sym}}^{p,\tilde{p}}(x,\tilde{x})=\langle x-\tilde{x},p-\tilde{p}\rangle$. If $h$ is differentiable with unique subgradients, we will drop the reference to $p,\tilde{p}$ and just write $D_{h,\text{sym}}(x,\tilde{x})$. Recall that, if $h$ is convex, $D_{h,\text{sym}}$ is always non-negative.
Suppose now that the potential $h:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable, that $\nabla h$ is $L_{\nabla h}$-Lipschitz-continuous and $\omega_{h}$-strongly convex. We consider two coupled Markov chains $(X_{n})_{n},(\tilde{X}_{n})_{n}$ defined by initial values $X_{0}\sim\mu_{0},\tilde{X}_{0}\sim\tilde{\mu}_{0}$ and
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I)\vspace{0.4em}\\ X^{n+1}=X^{n}-\tau\nabla h(X_{n})+\sqrt{2\tau}\xi^{n+1}\vspace{0.4em}\\ \tilde{X}^{n+1}=\tilde{X}^{n}-\tau\nabla h(\tilde{X}^{n})+\sqrt{2\tau}\xi^{n+1}\end{array}\right.$
@@ 4 Convergence of the Discrete Time Algorithm
If $h$ is $\omega_{h}$-strongly convex, we can use
$D_{h,\textrm{sym}}(X^{n},\tilde{X}^{n})\geq\frac{\omega_{h}L_{\nabla h}}{\omega_{h}+L_{\nabla h}}\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\frac{1}{\omega_{h}+L_{\nabla h}}\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2},$
@@
$\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2}\leq\left(1-\tau\frac{2\omega_{h}L_{\nabla h}}{\omega_{h}+L_{\nabla h}}\right)\lVert X^{n}-\tilde{X}^{n}\rVert^{2}+\tau\left(\tau-\frac{2}{\omega_{h}+L_{\nabla h}}\right)\lVert\nabla h(X^{n})-\nabla h(\tilde{X}^{n})\rVert^{2}.$
If the step size is chosen to be $\tau=2/(\omega_{h}+L_{\nabla h})$, we obtain the convergence rate
$\lVert X^{n}-\tilde{X}^{n}\rVert^{2}\leq\left(\frac{L_{\nabla h}-\omega_{h}}{L_{\nabla h}+\omega_{h}}\right)^{2n}\lVert X^{0}-\tilde{X}^{0}\rVert^{2}.$
Denoting the distribution of $X^{n}$ by $\mu_{n}$ and that of $\tilde{X}^{n}$ by $\tilde{\mu}_{n}$, we obtain by taking the infimum over all couplings of $\mu_{0}$ and $\tilde{\mu}_{0}$ on the right side that
$\mathcal{W}_{2}(\mu_{n},\tilde{\mu}_{n})\leq\left(\frac{L_{\nabla h}-\omega_{h}}{L_{\nabla h}+\omega_{h}}\right)^{n}\mathcal{W}_{2}(\mu_{0},\tilde{\mu}_{0}).$
This shows that ULA is stable and will converge to its unique stationary distribution. Similar computations can be carried out for other time-discretized schemes based on Langevin diffusion, e.g. time-implicit leading to proximal sampling schemes. Since the stationary distribution of the time-discrete scheme is not the target distribution $\mu^{\ast}$, the missing step to relate the samples drawn from the Markov chain and the target is to couple a time-discrete chain and a time-continuous process and then finding an error bound. Typically, this is more challenging and the introduced bias (in Wasserstein distance) between the two settings scales with a power of the step size $\tau$.
4.1 Stability of Primal-Dual Langevin Sampling
We now turn to the primal-dual setting of (1.4). Consider the following coupling in discrete time with initial values $(X^{0},Y^{0})\sim\pi_{0}$ and $(\tilde{X}^{0},\tilde{Y}^{0})\sim\tilde{\pi}_{0}$ and identical stochastic input to both iterations:
$\left\{\begin{array}[]{l}\xi^{n+1}\sim\mathrm{N}(0,I)\vspace{0.8em}\\ Y^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(Y^{n}+\sigma KX_{\theta}^{n})\\ X^{n+1}=\operatorname{prox}_{\tau g}(X^{n}-\tau K^{T}Y^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ X_{\theta}^{n+1}=X^{n+1}+\theta(X^{n+1}-X^{n})\vspace{0.8em}\\ \tilde{Y}^{n+1}=\operatorname{prox}_{\sigma f^{\ast}}(\tilde{Y}^{n}+\sigma K\tilde{X}_{\theta}^{n})\\ \tilde{X}^{n+1}=\operatorname{prox}_{\tau g}(\tilde{X}^{n}-\tau K^{T}\tilde{Y}^{n+1})+\sqrt{2\tau}\xi^{n+1}\\ \tilde{X}_{\theta}^{n+1}=\tilde{X}^{n+1}+\theta(\tilde{X}^{n+1}-\tilde{X}^{n})\end{array}\right.$ (4.1)